OpenStudy (anonymous):

Find all solutions of the equation and express them in the form a + bi. x^2 + x + 1 = 0

OpenStudy (anonymous):

(x+1/2)^2 + 3/4 = 0 (x+1/2)^2 = -3/4 \[x = \pm \sqrt{-3/4} -1/2\] \[x = -1/2 \pm \frac{\sqrt{3}}{2} * i\]